Consider an ice cube of edge 1.0 cm kept in a gravity free hall. Find the surface area of the water when the ice melts. Neglect the difference in densities of ice and water.

Grade:10

4 Answers

Deepak Patra
9 years ago
Sol. Surface area = 4πr2
Koustav Saha
9 years ago
Consider an ice cube of edge 1.0 cm kept in a gravity free hall.Find the surface area of water when the ice melts
 
Subodh Wasnik
6 years ago
I am unable to solve it. I tried very but cannot get the steps for it. However i know the answer. Answer is (9π)^1/3
Subodh Wasnik
6 years ago
Since ice is in gravity free space, thus after melting it forms a sphere. Area of ice cube= area of sphere1 cm3 =4/3πr^3=> r=(3/4π)^1/3Now, surfacr area is given byS.A.= 4πr^2 =4π(3/4π)^2/3 =(9π)^1/3